OpenStudy (anonymous):

if you save $4,500 at an interest rate of 13 percent per year how much will you have at the end of seven years?

OpenStudy (anonymous):

13% of $4,500 is $585 $585 x 7 =$4,095 c=

OpenStudy (anonymous):

so did i win the medal c=

OpenStudy (anonymous):

try future value=p*(1+i)t Present value=p/ (1+i)t

OpenStudy (anonymous):

open a now Question and can i get best response please

OpenStudy (anonymous):

i think it is 0 but i dont know sorry